# Retinomorphic Feature Detection and Machine Vision in a Network Laser

### v1

@@ Abstract
Biological vision systems seamlessly integrate feature-sensitive spatial processing and strong nonlinearity. The retina performs feature detection via ganglion cells, which nonlinearly enhance image features such as edges through lateral inhibition processes between neighbouring cells. As demand for specialised machine learning hardware increases, physical systems are sought which combine feature detection with the nonlinear processing required for tasks such as image classification. Currently, physical machine vision schemes detect features linearly, then employ separate digital systems for subsequent nonlinear activation.
Here, we introduce a bio-inspired ‘retinomorphic’ machine vision platform using a semiconductor network laser. The system hosts many spatially-overlapping lasing modes which detect multiple image features in parallel via their lasing amplitude. Integrated nonlinearity is provided by antagonistic competition for gain between modes - a photonic analogue of the lateral inhibition in retinal cells. Detected feature maps are fed back through the system, providing multi-layer image classification with intrinsic nonlinear processing. Accuracies of 98.05% and 87.85% are achieved for MNIST digits and Fashion-MNIST tasks respectively.
@@ Introduction
The drawback of such linear physical approaches is that they are unable to provide the crucial nonlinearity required for more complex machine vision tasks beyond feature detection, such as image classification. This is typically navigated by passing the detected feature maps to a digital CMOS computer which implements nonlinear activation in software[11, 12, 13]. Often, the software also performs the full classification process via a conventional deep neural network with additional nonlinear activations[8, 10]. This is nonetheless excellent progress - convolutional feature detection is computationally expensive in software, hence appealing to offload to specialised physical systems.
@@
It is appealing to draw inspiration from biological architectures, and engineer physical platforms capable of detecting multiple features in parallel with integrated nonlinearity. Neuromorphic computing is an approach which takes this bio-inspired angle and aims to implement machine learning by leveraging the complex nonlinear dynamics of physical systems[19, 11, 13]. However, gaining fine control over the nonlinear physical dynamics powering neuromorphic computing is a subtle and challenging task, particularly for more sophisticated multi-layer learning architectures such as CNNs where each layer of the neuromorphic system must perform very different processing[13, 12]. To date, a physical system integrating feature detection and physical nonlinearity remains unrealised to the best of our knowledge. Gaining enhanced control over nonlinear physical dynamics to realise such functional integration has been identified as a key goal for the field[11, 12, 13].
@@
Figure 1: Image feature detection in a complex network laser.
a)
Digital micromirror device takes 633 nm pump light and spatially structures it to an arbitrary input image.
b) Input image is illuminated onto 150 $\upmu$m InP network, triggering optical gain in the illuminated InP regions (highlighted in deeper red) and exciting photons which scatter through the network to produce many lasing modes at different wavelengths determined by the scattering paths. Spatial mode profiles of three lasing modes are indicated via different coloured nodes.
c) Lasing spectra emitted by the network are recorded via spatially-resolved spectrometer, giving ‘hyperspectral’ data with spectra recorded for one-dimensional slices across the network with approximately 1 $\upmu$m width. Modes detecting image features are highlighted by coloured boxes, corresponding to the feature maps in e) and f).
d) Input image is fractured into NxN pixel windows (here 4x4) and sequentially illuminated on the InP network as a raster scan. Small window on the person’s arm corresponds to the image illuminated in a-c).
e) Feature maps are produced by plotting the lasing amplitude of different modes. Each colour in the composite feature map corresponds to the lasing amplitude of a different mode, with 6 modes used to build this composite map.
f) Feature maps are shown for 6 separate modes, highlighting the range of image features detected in parallel by the system.
g) Examples of how different spatial illumination patterns produce different hyperspectral output. Coloured boxes indicate location of 10 feature-detecting modes.
h) Lasing spectra recorded under full illumination at different pump fluence. Lasing threshold behaviour is shown in inset.
@@
The physical system employed here is a 150 $\upmu$m diameter InP network laser comprising interconnected lateral nanoscale waveguides, lithographically-patterned on a Si chip[28, 29]. The network design is a random Voronoi topology with three waveguides meeting at each vertex (average waveguide length = 5 $\upmu$m). A pump beam, spatially-structured by a digital micromirror device (DMD) (figure 1a), induces optical gain in the InP network (figure 1b). Photons are created by spontaneous emission and scatter through the network of waveguides, amplified by stimulated emission from the illuminated InP. Constructive interference along specific scattering paths defines the optical modes, which undergo net gain or loss depending on the extent of overlap between the illumination pattern and the mode profiles.
The large number of possible scattering paths results in many spatially-distributed overlapping modes which nonlinearly compete for gain. ump illumination is supplied by a 633 nm, 200 fs pulse laser with 1 kHz repetition rate, with pulse energy measured at 43 nJ at the sample. The spatial distribution of light emitted from the network by three different modes are highlighted by the coloured nodes in figure 1b), with the data experimentally imaged by a frequency-resolved slit-scan camera. Fabrication and optical setup details are provided in the Methods and supplementary figure S1.
@@
Image feature maps are extracted by recording the amplitude of different modes as small ‘windows’ of an input image are scanned across the InP network (fig. 1d), here 4$\times$4 pixel windows. Modes effectively provide functionality akin to convolutional filters, with lasing intensity increased or suppressed by the presence of different image features due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. Specific modes lase strongly when a given image feature is present (e.g. a left or right hand edge), providing direct feature detection via the recorded spectrum. Examples of spectral output and mode intensities varying in response to different input images are shown in supplementary figure S2.
For a single input illumination, all modes undergo their lasing processes in parallel and multiple feature maps are obtained by plotting the lasing amplitude of selected modes at each position of the convolutional scan. A composite feature map comprising the lasing output of six modes is shown in figure 1e). The different colours on the composite map correspond to features detected by different spectral regions, indicated by boxes on figure 1c) (additional details in supplementary figure S2. Figure 1f) shows the six separate feature maps used to compose the composite map. As the modes providing the different feature maps overlap spatially within the same compact physical network, the scheme provides a space-efficient platform for parallel feature detection.
Figure 2: Parallel image feature-detection via network lasing modes. a-f) are simulation results, g-k) are experimental.
@@
Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network, modes detecting left edges are located towards the left of the network, and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary figure S3, alongside the full set of 172 feature maps from all modes in supplementary figure S4. A regular hexagonal network was simulated to compare with the random graph topology with far fewer detected features, shown in supplementary figure S5.
@@
Figure 3: Multilayer image classification.
a)
Feature detection: Input images are fractured into 4x4 pixel windows which are illuminated on the InP network in a sequential raster scan. The lasing amplitude of ten selected modes are used to perform spectrally-multiplexed feature detection and generate ten feature maps (three shown here, all ten shown in SI).
b) Classification: Detected feature maps are then illuminated onto the network as whole images in a single illumination step with no window scanning. Output spectra are recorded for each feature map, giving ten spectra per image.
c) The ten spectra are then combined in parallel, and a single logistic regression step is applied during training to generate a set of weights, one per hyperspectral channel (36813 channels after feature selection for MNIST-digits). For testing, no further regression is applied - each hyperspectral channel is multiplied by the fixed weights for each class obtained during training then summed. The class receiving the highest sum is then given as the prediction. A test accuracy of 98.05% is achieved for MNIST-digits.
d) The same architecture is employed for the Fashion-MNIST task. Two input images are shown with six feature maps, colours corresponding to the modes shown in figures 1 and 2.
e) A test accuracy of 87.85% is achieved for Fashion-MNIST.
f,g) Training length vs. test classification accuracy comparison for MNIST digits (f) and Fashion-MNIST (g), with a zoomed-in plot of classification performance at shorter training sets for each task. The photonic convolutional network is compared against four software approaches: Logistic regression performed on the raw image data, a multi-layer perceptron, a LeNet-5 software CNN and a larger CNN (network details in text).
@@ Conclusions
As the lasing modes responsible for feature detection spatially overlap, within the same InP network we are able to provide compact feature detection of multiple image features. In contrast, existing physical convolutional hardware schemes typically detect one feature per repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected. While we currently experimentally detect ten parallel image features, we find up to forty in simulation and the number of detected features is likely to increase as we deepen understanding of the links between network topology and feature sensitivity.
@@
As we develop finer control over the network dynamics, one can envision networks engineered to detect arbitrary specific features or provide reconfigurable detection via altering network dynamics via electrical gating or selective illumination of network sub-regions. More broadly, our approach shows the benefits of designing photonic computing systems with intrinsic physical nonlinearities and a dynamic balance of competing excitatory and inhibitory interactions.
@@
AF designed and produced the figures with contributions from JD.
@@ Methods
The lasing modes and intensities of the network are modelled using netSALT[24]. The designed network is represented as a graph, with the edges subdivided to overlap with an 8x8 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency $k_{a}=7$ with a width $\gamma_{a}=0.35$, and are pumped up to a normalised power $D_{0}=0.5$ (For a detailed physical explanation of the parameters, see Ref. 24). A separate simulation is performed for each 8x8 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane $k$, and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im$(k)=0$], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.
@@
Figure S2 f) shows an example of how varying the spatial pump illumination pattern results in different lasing output responses by the network. As the spatial illumination pattern in the ‘image windows’ column is varied, the hyperspectral response shifts (middle column) and the amplitude of the feature-detecting modes varies dynamically depending on which image features are present (right-hand column).
@@
f) Varying hyperspectral response and feature-detecting mode intensities as the input illumination pattern is varied, as when windows of an input image are illuminated on the network.

### v2

# Retinomorphic Machine Vision in a Network Laser

@@ Abstract
With the growing prevalence of AI, demand increases for efficient machine learning hardware. Physical systems are sought which combine image feature detection with the essential nonlinearity for tasks such as image classification. Existing physical hardware typically detects features linearly, then employs digital processing for nonlinear activation. Biological vision systems excel at nonlinear image processing. The retina detects features in ganglion cells via lateral inhibition, where neighbouring cells nonlinearly compete for activation.
We present a bio-inspired ‘retinomorphic’ machine vision platform using an on-chip semiconductor network laser. The system detects multiple features in parallel via spatially-overlapping lasing modes, with integrated nonlinearity provided by antagonistic gain competition between modes - a photonic analogue of retinal inhibition. Parallel feature-detection enhances efficiency relative to feature-detection schemes which operate sequentially or via multiple device copies, with Si-compatible processing and a compact $\upmu$m-scale footprint relative to existing mm-scale systems. We report 98.05% accuracy on MNIST-digits and 87.85% on Fashion-MNIST, with strong performance on short training datasets.
@@ Introduction
The drawback of such linear physical approaches is that they are unable to provide the crucial nonlinearity required for more complex machine vision tasks beyond feature detection, such as image classification. This is typically navigated by passing the detected feature maps to a digital CMOS computer which implements nonlinear activation in software[11, 12, 13]. Often, the software also performs the full classification process via a conventional deep neural network with additional nonlinear activations[8, 10]. A multilayer optical scheme has recently been demonstrated using saturating opto-electronic nonlinearity between two linear optical layers, employing a fully-connected rather than convolutional architecture[14].
@@
It is appealing to draw inspiration from biological architectures, and engineer physical platforms capable of detecting multiple features in parallel with integrated nonlinearity. Neuromorphic computing is an approach which takes this bio-inspired angle and aims to implement machine learning by leveraging the complex nonlinear dynamics of physical systems[20, 11, 13]. However, gaining fine control over the nonlinear physical dynamics powering neuromorphic computing is a subtle and challenging task, particularly for more sophisticated multi-layer learning architectures such as CNNs where each layer of the neuromorphic system must perform very different processing[13, 12]. To date, a physical system integrating convolutional feature detection and physical nonlinearity remains unrealised to the best of our knowledge. Gaining enhanced control over nonlinear physical dynamics to realise such functional integration has been identified as a key goal for the field[11, 12, 13].
@@
Figure 1: Image feature detection in a complex network laser. a) Digital micromirror device spatially structures 633 nm pump light into an arbitrary input image, which is illuminated onto a 150 $\upmu$m InP network b). The pump light triggers optical gain in illuminated regions of InP network, exciting photons which scatter through the network to produce many lasing modes at different wavelengths determined by the scattering paths. Spatial mode profiles of three modes are indicated via different coloured nodes. c) Lasing spectra emitted by the network are recorded via spatially-resolved spectrometer, giving ‘hyperspectral’ data with spectral and spatial dimensions , at $\sim$1 $\upmu$m spatial resolution. Spectral regions which detect image features are highlighted by coloured boxes, corresponding to feature maps in i) and j). d) Mode intensity vs. pump power for all modes, with different thresholds and nonlinear behaviours seen between modes. Colour-coded modes correspond to the colour-coded spectral regions and feature maps in i,j). e) Schematic of feature detection via kernel filters in software convolutional neural networks. Kernel filters are raster-scanned over the input image in small ’windows’ performing sequential multiply-and-accumulate operations to produce feature maps. Each image feature (e.g. left or right edges) is detected by a separate kernel filter, detecting multiple features consumes more time and energy. The process is linear, so subsequent nonlinear activations must be performed on the feature maps to enable tasks such as image classification. f) Schematic of network laser feature detection. Small input image windows are sequentially illuminated on the InP network. Different network lasing modes detect multiple image features in parallel from a single illumination, with integrated physical nonlinearity. Experimentally, ten features are concurrently detected with forty detected in simulation. The network laser efficiently integrates multiple steps of the feature detection process, which must be performed sequentially in software. g) Input ‘cameraman’ image, inset shows example image windows (blue boxes) and raster-scan direction. h) Image windows from g) with spectral output of the InP network and mode intensities for six feature-detecting lasing modes (colours follow spectral regions and feature maps in c,i,j). Mode amplitudes vary in response to the presence of different image features. i) Composite feature map detected by the InP network. Different feature colours correspond to the lasing amplitude of the six lasing modes indicated in c,h,j). A range of image features are detected in parallel. j) Feature maps for the six separate lasing modes which form the composite map, highlighting the range of parallel features detected by the system.
@@
Figure 1e) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster scanned across small image ‘windows’ of an input image, performing multiply-and-accumulate operations via matrix multiplication. The kernel values are multipled against the image pixel values, the results summed to a single number which becomes the pixel value of an output feature map at the centre of the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming additional time and energy for each feature detected. This process is entirely linear, so subsequent nonlinear activations must be performed at time and energy cost to use the feature maps for ML tasks such as image classification. This linear approach is what has so-far been used for physical convolutional feature detection schemes[8, 9, 7, 10].
Figure 1f) shows a schematic of the network laser scheme. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, but here the network simultaneously detects many features in parallel via multiple lasing modes, with integrated physical nonlinearity. This increases efficiency relative to software CNNs and their physical counterparts, with the network laser currently detecting ten parallel features experimentally (forty in simulation) in a compact 150 $\upmu$m footprint, relative to the mm-scale of alternative hardware schemes.
Figure 1g-j) show experimental parallel feature detection for a ‘cameraman’ test image (g). Example image windows ($\times 4$ pixels, inset of g) are shown in (h) with their respective network lasing spectra and mode intensities for six feature-detecting modes. Mode lasing intensity is increased or suppressed by the presence of different image features due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. Specific modes lase strongly when a given image feature is present (e.g. a left or right hand edge), providing direct feature detection via the lasing amplitude. This enables feature maps to be obtained by plotting the lasing amplitude of different modes at different positions in the raster-scan. Figure 1i) shows a composite feature map produced by six lasing modes, with the different colours corresponding to the lasing amplitude of the six modes indicated in a,e,g). A broad range of features are detected in parallel, with separate output for each feature provided by the different spatial and spectral locations of the modes. Figure 1j) shows separate feature maps for the six modes comprising the composite map. Feature maps for all ten feature-detecting modes are shown in supplementary figure S10.
Figure 2: Parallel image feature-detection via network lasing modes. a-f) show simulation results, g-k) are experimental.
@@
Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network, modes detecting left edges are located towards the left of the network, and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary figure S3, alongside the full set of 172 feature maps from all modes in supplementary figure S4. A regular hexagonal network was simulated to compare against the random graph topology (supplementary figure S5), with far fewer detected features in the regular network showing the benefits of a disordered graph topology. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance fruitful grounds for future work.
@@
Figure 3: Multilayer image classification. a) Feature detection: Input images are fractured into 4x4 pixel windows which are illuminated on the InP network in a sequential raster scan. The lasing amplitude of ten selected modes are used to perform spectrally-multiplexed feature detection and generate ten feature maps (three shown here, all ten shown in SI). b) Classification: Detected feature maps are then illuminated onto the network as whole images in a single illumination step with no window scanning. Output spectra are recorded for each feature map, giving ten spectra per image. c) The ten spectra are then combined in parallel, and a single logistic regression step is applied during training to generate a set of weights, one per hyperspectral channel (36813 channels after feature selection for MNIST-digits). For testing, no further regression is applied - each hyperspectral channel is multiplied by the fixed weights for each class obtained during training then summed. The class receiving the highest sum is then given as the prediction. A test accuracy of 98.05% is achieved for MNIST-digits. d) The same architecture is employed for the Fashion-MNIST task. Two input images are shown with six feature maps, colours corresponding to the modes shown in figures 1 and 2. e) A test accuracy of 87.85% is achieved for Fashion-MNIST. f,g) Training length vs. test classification accuracy comparison for MNIST digits (f) and Fashion-MNIST (g), with a zoomed-in plot of classification performance at shorter training sets for each task. The photonic convolutional network is compared against four software approaches: Logistic regression performed on the raw image data, a multi-layer perceptron, a LeNet-5 software CNN and a larger CNN (network details in text).
@@ Conclusions
As the lasing modes responsible for feature detection spatially overlap, within the same InP network we are able to provide compact feature detection of multiple image features. In contrast, existing physical convolutional hardware schemes typically detect one feature per repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected. While we currently experimentally detect ten parallel image features, we find up to forty in simulation and the number of detected features is projected to increase as we deepen understanding of the links between network topology and feature sensitivity.
@@
As we develop finer control over the network dynamics, one can envision networks engineered to detect arbitrary specific features or provide reconfigurable detection via altering network dynamics via electrical gating or selective illumination of network sub-regions. More broadly, our approach shows the benefits of designing physical computing systems with intrinsic physical nonlinearities and a dynamic balance of competing excitatory and inhibitory interactions.
@@
AF designed and produced the figures with contributions from JD and WKN.
@@ Methods
The lasing modes and intensities of the network are modelled using netSALT[25]. The designed network is represented as a graph, with the edges subdivided to overlap with an 8x8 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency $k_{a}=7$ with a width $\gamma_{a}=0.35$, and are pumped up to a normalised power $D_{0}=0.5$ (For a detailed physical explanation of the parameters, see Ref. 25). A separate simulation is performed for each 8x8 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane $k$, and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im$(k)=0$], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.